A new computer virus can enter the system through e-mail or through the internet. There
is a 30% chance of receiving this virus through e-mail. There is a 40% chance of
receiving it through the internet. Also, the virus enters the system simultaneously
through e-mail and the internet with probability 0.15. What is the probability that the
virus does not enter the system at all?

Answer by ikleyn(53856)

.

The probability that the virus will enter the computer via Email OR the Internet is


    p = 0.3 + 0.4 - 0.15 = 0.55.


The probability that the virus will not enter the computer at all is the complement to it, 


which is  P' = 1 - P = 1 - 0.55 = 0.45.      ANSWER

Solved.